# Online Vectorized HD Map Construction using Geometry

### v1

@@ Abstract
The construction of online vectorized High-Definition (HD) maps is critical for downstream prediction and planning. Recent efforts have built strong baselines for this task, however, shapes and relations of instances in urban road systems are still under-explored, such as parallelism, perpendicular, or rectangle-shape. In our work, we propose GeMap (Geometry Map), which end-to-end learns Euclidean shapes and relations of map instances beyond basic perception. Specifically, we design a geometric loss based on angle and distance clues, which is robust to rigid transformations. We also decouple self-attention to independently handle Euclidean shapes and relations. Our method achieves new state-of-the-art performance on the NuScenes and Argoverse 2 datasets. Remarkably, it reaches a 71.8% mAP on the large-scale Argoverse 2 dataset, outperforming MapTR V2 by +4.4% and surpassing the 70% mAP threshold for the first time. Code is available at https://github.com/cnzzx/GeMap.
@@ 1 Introduction
Vectorized HD maps provide sparse and structured environmental information for autonomous vehicles and have been widely adopted in downstream tasks, such as trajectory forecasting [17, 3, 46] and planning [32, 5]. Online Vectorized HD Map Construction can significantly reduce the need for labor-intensive annotations and facilitate real-time updates in autonomous driving [18, 4, 14].
@@
Figure 1: Invariant Geometric Properties. As the ego vehicle moves, after rotation ($\leftarrow$) or translation ($\rightarrow$), the shape of this crossing and relationship (perpendicular) between other lanes stays unchanged, which indicates the invariant property of geometry to rigid transformations.
@@
With the development of Bird’s-Eye-View (BEV) representation, online HD map construction has achieved significant advancements [19, 16, 9]. Early works [11, 27, 30, 42] formulate HD Map construction as a dense prediction task. However, these methods generate maps in image format, which is redundant for representing sparse map instances. Then, a more compact map formulation was introduced to minimize redundancy, albeit at the cost of adding time-consuming post-processing steps [14]. In response to this, recent works [21, 18, 31, 4] attempt to construct vectorized HD maps end-to-end to avoid extensive post-processing. These methods typically sample points from map instances and represent each instance as a polyline [21, 18, 4] or parameterized curve [31].
@@
We observe that urban road systems exhibit significant geometric characteristics, such as parallel lanes, perpendicular crossings, equal lane widths, etc. However, these geometric properties of shapes and relations between map instances have not been fully explored. Meanwhile, there are two notable limitations among existing methods [21, 18, 31, 4] which can be alleviated by leveraging these geometric properties. 1) As the ego-vehicle moves, instances experience rotations and translations, as depicted in Figure 1. However, the polylines and parameterized curve representations employed in prior research, which rely on the absolute coordinates of points, are inherently sensitive to these rotational and translational changes [7, 31]. 2) Geometric properties widely exist in driving scenarios such as rectangle shape, parallelism, and perpendicular relationships, as illustrated in Figure LABEL:fig:idea-a. Incorporating these properties could significantly enhance HD map construction. However, the vanilla attention mechanism struggles to effectively learn diverse shape and relation geometry due to objective conflicts, as detailed in § 3.5.
Figure 2: Geometric properties and G-Representation. (a) Geometry in urban road system. For example, pedestrian crossings are near rectangular intersections and are always vertical to road boundaries. (b) (c) We propose to utilize angles and distances to model local geometry in a map instance and between different instances, denoted as Euclidean Shape Clues (§ 3.3.1) and Euclidean Relation Clues(§ 3.3.2), respectively. Note that included angles and distances between map instances are invariant to positional and directional changes of an ego vehicle.
@@
To address the above two limitations, we introduce a geometric representation that captures shapes of individual map instances and relations between different instances as illustrated in Figure LABEL:fig:idea-b and LABEL:fig:idea-c, referred to as G-Representation. It enhances the vanilla representation of map instances by incorporating a translation- and rotation-invariant representation that effectively leverages instance geometry. The local structures of map features are encoded using displacement vectors. These vectors are computed from the absolute coordinates of polyline points to effectively represent the relative positions and orientations of adjacent points. To quantify these features within Euclidean space, we employ both the magnitudes of the displacement vectors and the angles formed between these vectors. Based on G-Representation, we propose Euclidean Shape Clues to represent the shapes of individual map instances and Euclidean Relation Clues to represent the inter-instance relations in Euclidean space.
@@
G-Representation addresses the abovementioned two limitations in an elegant way. 1) By concentrating on the relative relationships between points within and across map instances, it inherently attains translation and rotation invariance. This enhances its robustness against variations in data collection and equips it to effectively handle different coordinate systems. 2) It captures the inherent geometry in urban road systems. For instance, in Euclidean space, the geometric representation of parallel lanes is indicated by an almost negligible angle between their displacement vectors, approaching zero.
@@
Building upon G-Representation, we propose a framework named GeMap for HD map construction. It includes a BEV encoder to extract BEV features from multi-view input images and a geometry-decoupled decoder with a refined focus on geometric aspects. Specifically, we adapt the attention mechanism [37] and propose Geometry-Decoupled Attention (GDA). GDA sequentially applies attention to queries belonging to the same instance and attention to queries across different instances. This can significantly boost the learning of key points of shapes and relations between map instances. Furthermore, we propose an objective function named Euclidean Loss to optimize G-Representation. Specifically, we transform the conventional polylines of the ground truth map into our G-Representation. In this way, the model gets optimized to better understand the magnitude of displacement vectors and the angles between them, thereby facilitating a more effective learning of geometric properties.
@@
Experiments on NuScenes and Argoverse 2 datasets demonstrate the effectiveness of GeMap. We reach new state-of-the-art performances on both datasets. With camera images only, GeMap achieves 69.4% and 71.8% mAP on the Nuscenes and Argovserse 2, respectively.
Our contributions are summarized as follows.
@@
  • We propose G-Representation which harnesses critical geometric properties of rotational and translational invariance in autonomous driving scenes, opening up new research avenues within the field.
@@
  • We introduce GeMap, an innovative framework incorporating geometry-decoupled attention and Euclidean Loss function, specifically designed to enhance geometric learning capabilities in Online HD Map.
@@ 2.2 Cross-view BEV Learning
In autonomous driving systems, the conversion of Perspective View (PV) camera images into a unified BEV space is a significant challenge. While previous studies such as [30, 15] use depth estimation from monocular images for this transformation, others [2, 16, 23] have developed PV-to-BEV conversion methods without explicit depth details. For example, GKT [2] employs a geometric-guided kernel, and BEVFormer [16] uses deformable attention for the BEV conversion. However, these methods can distort shapes and geometric relations, leading to inaccuracies. Our approach integrates geometric supervision into prediction, addressing these geometric inconsistencies while acknowledging the value of depth information.
@@ 2.3 Geometric Instance Modeling
HD maps, depicting diverse geometric features such as pedestrian crossings and lane boundaries, present vectorization challenges. Traditional vectorization approaches model these features as polylines [33, 38, 21, 18] or polynomial curves [36, 34, 6, 31], with some incorporating Bézier Curves for improved fitting [6, 31]. However, such methods often overlook geometric properties including shapes, parallelism, etc. Research on geometric loss has mainly addressed rasterization-based shape [36, 44] or edge loss [18], and fails to preserve cross-instance relationships. PivotNet [4] offers an architectural advancement with its line-aware point decoder, yet it still does not fully address the complexity of instance geometry.
Given the current under-exploration of diverse geometries in shape and relation, we propose utilizing Euclidean shape and relation losses and decoupling the traditional self-attention module to empower the model with a more robust understanding of instance geometry.
@@ 3.1 Preliminary
The architecture of our method is illustrated in Figure 3. Input images are denoted by ${\mathbb{I}}=\{{\bm{I}}_{i}\}_{i=1}^{N_{c}}$, where ${\bm{I}}_{i}\in\mathbb{R}^{H\times W\times 3}$ and $N_{c}$ is the sample number. Output is a set of $N$ map instances ${\mathbb{M}}=\{{\bm{L}}_{i}\}_{i=1}^{N}$, where each map instance is represented by a polyline ${\bm{L}}_{i}\in\mathbb{R}^{N_{v}\times 2}$, i.e., an ordered sequence of $N_{v}$ two-dimensional points. As shown in Table 3, the commonly adopted pipeline comprises three steps. 1) A BEV feature extractor processes the multi-view images. 2) The extracted BEV features are fed into a map decoder, which predicts the map instances based on the BEV features. 3) To optimize the pipeline, the predicted map instances are compared against the ground truth. Beyond point-to-point comparison, G-representations of predictions and the ground truth are computed and the difference is measured by $L^{1}$ loss.
@@ 3.2 Architecture Overview
Geometric Loss. For training, we convert both the predicted polylines and ground truth into the proposed G-Representation. Based on that, we let the model optimize the predicted lengths of displacement vectors and angles between displacement vectors to match the ground truth.
@@ 3.3 Geometric Representation
3.3.1 Euclidean Shape Clues
We first introduce the representation of shapes of individual map instances. For each instance, we describe the local geometry with displacement vectors between neighboring points. The displacement vectors are computed as
@@ 3.3.1 Euclidean Shape Clues
Figure 4: Euclidean shape clues. Lengths of displacement vectors and angles between neighboring vectors consist of shape clues and are utilized to compute shape loss. The right part shows how to connect Euclidean shape clues to shape geometry.
These displacement vectors are sufficient to represent the shape of a map instance and such a representation is invariant to rigid transformations. However, we would like to note that this representation is vulnerable to rotations, which might prevent the model from learning robust geometry. To solve this problem, we propose to represent the shape with lengths of displacement vectors and angles between consecutive displacement vectors, as illustrated in Figure 4. Specifically, the $i$-th instance is denoted by $N_{v}$ angle values and $N_{v}$ distance values. Let $u$ be the index, the $u$-th angle value and distance value of the $i$-th instance are computed as
@@
The proposed representation inherently captures common geometric patterns, such as parallelism, right angles, and proper line width, by translating them into corresponding numerical patterns within this representation. For example (Figure 4), a rectangle is characterized by one $90^{\circ}$ angle and equal lengths of opposite displacement vectors; in addition, a straight line is represented by a $0^{\circ}$ angle.
3.3.2 Euclidean Relation Clues
Having highlighted the translation and rotation invariance of G-Representation and its advantages in representing the geometric shapes of individual map instances, we further introduce its ability to capture the relations between two map instances, e.g., parallelism and perpendicularity between two instances in the Euclidean space.
Figure 5: Euclidean relation clues. Distances between pairs of points and angles between pairs of displacement vectors indicate relation clues. Such relation clues are more superficially connected to Euclidean relation geometry as shown in the right part.
Specifically, given the vanilla representations and displacement vectors of the $i$-th and $j$-th map instance, we represent the relation between instance $i$ and $j$ with 1) angles between each pair of their respective displacement vectors, and 2) distances between each pair of their points, as illustrated in Figure 5. Formally,
@@ 3.3.2 Euclidean Relation Clues
This representation is also translation- and rotation-invariant. Similar to Euclidean Shape Clues, it inherently captures common relations between map instances by translating them into corresponding numerical patterns. For example, Figure 5 shows that the perpendicular relation can be directly represented by a $90^{\circ}$ angle, and the distance between two parallel polylines, which may correspond to the width of a lane, is naturally represented by point-to-point distances.
@@ 3.4 Euclidean Loss and Objectives
We transform both the ground truth data and the model’s predicted polylines from their original format into G-Representation. This conversion allows the model to independently optimize the angles between displacement vectors and the distances between points to more accurately align with the ground truth. The proposed Euclidean Loss is composed of two parts that measure how accurately the model predicts the shape of individual map instances and the inter-instance relations, respectively, which are denoted by $\mathcal{L}_{\rm shp}$ and $\mathcal{L}_{\rm rel}$. The Euclidean Loss can be computed as the following:
$\mathcal{L}_{\rm Euc}=\lambda_{1}\cdot\mathcal{L}_{\rm shp}+\lambda_{2}\cdot\mathcal{L}_{\rm rel}.$ (4)
Specifically,
$\mathcal{L}_{\rm shp}$ $=\sum_{i=1}^{N}\sum_{j=1}^{N_{v}}|\hat{d}^{i}_{j}-d^{i}_{j}|+\ell(\hat{\alpha}^{i}_{j},\alpha^{i}_{j})\,,$ (5)
$\mathcal{L}_{\rm rel}$ $=\sum_{i=1}^{N}\sum_{j>i}^{N}\sum_{u=1}^{N_{v}}|\hat{d}^{i,j}_{u}-d^{i,j}_{u}|+\sum_{v=1}^{N_{v}}\ell(\hat{\alpha}^{i,j}_{u,v},\alpha^{i,j}_{u,v})\,,$
where $\ell(\cdot)$ is a special loss function based on $L^{1}$ loss. We avoid using the inverse trigonometric function to directly compute the angles, but use sine and cosine values instead:
@@
Following the common practice [18, 43], we also use focal loss [20] for classification denoted by $\mathcal{L}_{\rm cls}$. For polyline regression, we adopt point-to-point loss and edge direction loss, denoted by $\mathcal{L}_{\rm pts}$ and $\mathcal{L}_{\rm dir}$ respectively. Following [31, 19], we further adopt segmentation-based loss $\mathcal{L}_{\rm seg}$ and depth estimation loss $\mathcal{L}_{\rm dep}$ (Details in Appendix). The overall loss function can be written as:
$+\beta_{3}\cdot\mathcal{L}_{\rm dir}+\beta_{4}\cdot\mathcal{L}_{\rm seg}+\beta_{5}\cdot\mathcal{L}_{\rm dep}.$
We would like to note that $\mathcal{L}_{\rm seg}$ and $\mathcal{L}_{\rm dep}$ are optional. We also experiment with a simpler objective function to evaluate the effect of our main contribution alone, rather than compete with the recent powerful baselines. That is,
$\mathcal{L}^{\prime}=\lambda\cdot\mathcal{L}_{\rm geo}+\beta_{1}\cdot\mathcal{L}_{\rm cls}+\beta_{2}\cdot\mathcal{L}_{\rm pts}+\beta_{3}\cdot\mathcal{L}_{\rm dir}.$ (8)
The most relevant hyperparameters to our contributions include $\lambda$, $\lambda_{1}$, and $\lambda_{2}$, whose effects on the model performances are evaluated in § 4.5.
Table 1: Comparison on the NuScenes dataset and our method reaches a new state-of-the-art performance. ”EB0”, ”R50”, ”PP”, ”Sec”, denote EfficientNet-B0 [35], ResNet50 [10], PointPillars [12], SECOND [41] respectively. Methods with two backbones utilize both camera and LiDAR inputs. “Seg. Loss” denotes whether any segmentation-based loss is adopted. The best result is highlighted in bold. We reproduce different methods on a single RTX3090 GPU to test FPS for fair comparison.
[…]

### v2

# https://invictus717.github.io/GeMap/

@@ Abstract
Online vectorized High-Definition (HD) map construction is critical for downstream prediction and planning. Recent efforts have built strong baselines for this task, however, geometric shapes and relations of instances in road systems are still under-explored, such as parallelism, perpendicular, rectangle-shape, etc. In our work, we propose GeMap (Geometry Map), which end-to-end learns Euclidean shapes and relations of map instances beyond fundamental perception. Specifically, we design a geometric loss based on angle and magnitude clues, robust to rigid transformations of driving scenarios. To address the limitations of the vanilla attention mechanism in learning geometry, we propose to decouple self-attention to handle Euclidean shapes and relations independently. GeMap achieves new state-of-the-art performance on the nuScenes and Argoverse 2 datasets. Remarkably, it reaches a 71.8% mAP on the large-scale Argoverse 2 dataset, outperforming MapTRv2 by +4.4% and surpassing the 70% mAP threshold for the first time. Code is available at https://github.com/cnzzx/GeMap.
Keywords:
HD Map Construction Geometry Representation Geometry-Decoupled Attention
@@ 1 Introduction
Vectorized HD maps provide structured environmental information for autonomous vehicles and have been widely adopted in downstream tasks, such as trajectory forecasting [19, 3, 49] and planning [35, 5]. Online Vectorized HD Map Construction can significantly reduce the need for labor-intensive annotations and facilitate real-time updates in autonomous driving [20, 4, 16].
@@
Figure 1: Geometric Invariance. (a) As the ego vehicle moves, after rotation ($\leftarrow$) and translation ($\rightarrow$), the shape of the crossing and the parallelism between lanes remain unchanged, which indicates the invariant property of geometry to rigid transformations. (b) Absolute coordinates are vulnerable to rotation and translation, however, our G-representation is invariant, which is more suitable to capture geometric properties.
@@
With the development of Bird’s-Eye-View (BEV) representation, online HD map construction has achieved significant advancements [21, 18, 9]. Early works [12, 30, 33, 45] formulate HD map construction as a dense prediction task. However, these methods generate maps in image format, which is redundant for representing sparse map instances. Then, a more compact map formulation was introduced to minimize redundancy, albeit at the cost of adding time-consuming post-processing steps [16]. In response to this, recent works [24, 20, 34, 4] attempt to end-to-end construct vectorized HD maps to avoid extensive post-processing. These methods typically sample points from map instances and represent each instance as a polyline [24, 20, 4] or parameterized curve [34].
@@
We observe that transportation road systems exhibit significant geometric characteristics (Figure LABEL:fig:idea-a), such as parallel lanes, perpendicular crossings, equal lane widths, etc. However, these geometric properties of shapes and relations between map instances have not been fully explored. Meanwhile, there are two notable limitations among existing methods [24, 20, 34, 4] which can be alleviated by leveraging these geometric properties. 1) Excessive dependency on absolute coordinates: as the ego-vehicle moves, instances experience rotations and translations, as depicted in Figure LABEL:fig:inv-a. However, widely adopted representations such as polylines and parameterized curves  [7, 34] are inherently sensitive to rotation and translation changes. 2) Objective conflicts of vanilla attention mechanism (detail in § 3.5) struggles to learn diverse shapes and relations, even though geometric properties such as rectangle shape, parallelism, and perpendicular relationships are commonly found in driving scenarios, as illustrated in Figure LABEL:fig:idea-a. We believe that incorporating these geometric properties significantly enhances the precision and efficiency of online HD map construction.
Figure 2: Geometric properties and G-Representation. (a) Geometry in the transportation road system. (b) (c) We propose to model the geometric properties of a single map instance and multiple instances, with magnitude $d$ and angle $\alpha$.
@@
To address the above two limitations, we introduce a novel geometric representation that captures shapes of individual map instances and relations between different instances as illustrated in Figure LABEL:fig:idea-b and LABEL:fig:idea-c, referred to as G-Representation. It enhances the vanilla representation of map instances by incorporating a translation- and rotation-invariant representation that effectively leverages instance geometry. The local structures of map features are encoded using displacement vectors. These vectors are computed from the absolute coordinates of polyline points to effectively represent the relative positions and orientations of adjacent points. To quantify these features within Euclidean space, we employ both the magnitudes of the displacement vectors and the angles formed between these vectors. Based on G-Representation, we propose the Euclidean Shape Clues to represent the shapes of individual map instances and the Euclidean Relation Clues to represent the inter-instance relations in Euclidean space.
@@
G-Representation is a simple yet effective method to address the abovementioned two limitations: 1) Geometric Invariance, by concentrating on the relative relationships between points within and across map instances, it inherently attains translation and rotation invariance as illustrated in Figure LABEL:fig:inv-b. This enhances its robustness against variations in data collection and equips it to effectively handle different coordinate systems. 2) Euclidean Modeling, it captures the inherent geometry in transportation road systems. Diverse shapes and relation geometry can be simplified as enumerations of magnitudes and directions. For instance, in Euclidean space, parallel lanes can be easily modeled as approximate directions (${\alpha\approx 0}$).
@@
Building upon G-Representation, we propose a framework named GeMap for HD map construction. A BEV encoder is used to extract features from multi-view input images, while a geometry-decoupled decoder is employed to focus on geometric aspects. Specifically, we adapt the attention mechanism [40] and propose Geometry-Decoupled Attention (GDA). GDA sequentially applies attention to queries belonging to the same instance and attention to queries across different instances. This can significantly boost the geometry learning of key points of shapes and relations between map instances. Furthermore, we propose an objective function named Euclidean Loss to optimize G-Representation. Specifically, we transform the conventional polylines of the ground truth map into our G-Representation. In this way, the model gets optimized to better understand the magnitude of displacement vectors and the angles between them, thereby facilitating a more effective learning of geometric properties.
@@
Experiments on nuScenes and Argoverse 2 datasets demonstrate the effectiveness of GeMap. We reach new state-of-the-art performances on both datasets. With camera images only, GeMap achieves 69.4% and 71.8% mAP on the nuScenes and Argovserse 2, respectively. Visualization results (§ 4.4) further demonstrate the better perception of shape and relation, alleviation of occlusion, and robustness to rigid transformations of GeMap.
Our contributions are summarized as the following:
@@
  • We propose G-Representation which harnesses critical geometric properties of rotation and translation invariance in autonomous driving scenarios, opening up new research avenues within the field.
@@
  • We introduce GeMap, a novel framework incorporating geometry-decoupled attention and Euclidean Loss function, specifically designed to learn the intrinsic geometry of online HD maps.
@@ 2.2 Cross-view BEV Learning
The conversion of Perspective View (PV) camera images into a unified BEV space is a significant challenge for autonomous driving systems. Previous studies such as [33, 17] use depth estimation from monocular images for this transformation, others [2, 18, 26] have developed PV-to-BEV conversion methods without explicit depths. For example, GKT [2] employs a geometric-guided kernel, and BEVFormer [18] uses deformable attention for the conversion. However, these methods can distort shapes and geometric relations, leading to inaccuracies. Our approach integrates geometric supervision in prediction, addressing these geometric inconsistencies while acknowledging the value of depth information.
@@ 2.3 Geometric Instance Modeling
HD maps depict instances with diverse geometric properties such as pedestrian crossings and lane boundaries, which present vectorization challenges. Traditional vectorization approaches model these instances as polylines [36, 41, 24, 20] or polynomial curves [39, 37, 6, 34], with some incorporating Bézier Curves for improved fitting [6, 34]. However, such methods often overlook geometric properties including shapes, parallelism, perpendicular, and etc. In 3D Lane Detection, some works [23, 11, 15] also explore simple geometry priors of lanes such as “equal lane width”. However, they are not generalizable to more complex HD map instances. Existing works on modeling geometry properties of map instances mainly address rasterization-based shape loss [39, 47] or edge loss [20], and they fail to preserve cross-instance relationships. PivotNet [4] offers an architectural advancement with its line-aware point decoder, yet it still does not fully address the complexity of instance geometry.
Given the current under-exploration of diverse geometric properties in shape and relation, we propose utilizing Euclidean shape and relation losses and decoupling the traditional self-attention module to empower the model with a more robust understanding of instance geometry.
@@ 3.1 Preliminary
The architecture of our method is illustrated in Figure 3. Input images are denoted by ${\mathbb{I}}=\{{\bm{I}}_{i}\}_{i=1}^{N_{c}}$, where ${\bm{I}}_{i}\in\mathbb{R}^{H\times W\times 3}$ and $N_{c}$ is the number of cameras. The output is a set of $N$ map instances ${\mathbb{M}}=\{{\bm{L}}_{i}\}_{i=1}^{N}$, where each map instance is represented by a polyline ${\bm{L}}_{i}\in\mathbb{R}^{N_{v}\times 2}$, i.e., an ordered sequence of $N_{v}$ two-dimensional points. As shown in Figure 3, the commonly adopted pipeline comprises three steps. 1) A BEV feature extractor processes the multi-view images. 2) The extracted BEV features are fed into a map decoder, which predicts the map instances based on the BEV features. 3) To optimize the pipeline, the predicted map instances are compared against the ground truth. Beyond point-to-point comparison, G-representations of predictions and the ground truth are computed and the difference is measured by $L^{1}$ loss.
@@ 3.2 Architecture Overview
Geometric Loss. For training, we convert both the predicted polylines and ground truth into the proposed G-Representation. Based on that, we let the model optimize the predicted magnitudes of displacement vectors and angles between displacement vectors to match the ground truth.
@@ 3.3 Geometric Representation
3.3.1 Euclidean Shape Clues.
We first introduce the representation of shapes of individual map instances. For each instance, we describe the local geometry with displacement vectors between neighboring points which are computed as:
@@
Figure 4: Euclidean Shape Clues. Magnitudes of displacement vectors and angles between neighboring vectors indicate shape clues and are utilized to compute shape loss. The right part shows how to connect Euclidean Shape Clues to shape geometry.
These displacement vectors are sufficient to represent the shape of a map instance and such a representation is invariant to translation transformations. However, we would like to note that this representation is vulnerable to rotations, which might prevent the model from learning robust geometry. To solve this problem, we propose to represent the shape with magnitudes of displacement vectors and angles between consecutive displacement vectors, as illustrated in Figure 4. Specifically, the $i$-th instance is denoted by $N_{v}$ angle values and $N_{v}$ magnitude values. Let $u$ be the index, the $u$-th angle value and the magnitude value of the $i$-th instance are computed as:
@@
The proposed representation inherently captures common geometric patterns, such as parallelism, right angles, and proper line width, by translating them into corresponding numerical patterns within this representation. For example, in Figure 4, a rectangle is characterized by one $90^{\circ}$ angle and equal magnitudes of opposite displacement vectors; in addition, a straight line is represented by $0^{\circ}$ angles. Beyond discussed regular cases, the Euclidean shape clues are capable of handling more complex shapes, as illustrated in scenarios (a) and (c) of Figure 6 and discussed in § 4.4.
3.3.2 Euclidean Relation Clues.
Having highlighted the translation and rotation invariance of G-Representation and its advantages in representing the geometric shapes of individual map instances, we further introduce its ability to capture the relations between two map instances, e.g., parallelism and perpendicular, in Euclidean space.
Figure 5: Euclidean Relation Clues. Angles between pairs of displacement vectors on different polylines, and magnitudes of displacement vectors between point pairs indicate relation clues. Such relation clues are more superficially connected to Euclidean relation geometry as shown in the boxes.
Specifically, given the vanilla representations and displacement vectors of the $i$-th and $j$-th map instance, we represent the relation between instance $i$ and $j$ with 1) angles between each pair of their respective displacement vectors, and 2) magnitudes of displacement vectors between each pair of points, as illustrated in Figure 5. Formally,
@@
This representation is also translation- and rotation-invariant. Similar to Euclidean Shape Clues, it inherently captures common relations between map instances by translating them into corresponding numerical patterns. For example, Figure 5 shows that the perpendicular relation can be directly represented by a $90^{\circ}$ angle, and the distance between two parallel polylines, which may correspond to the width of a lane, is naturally represented by magnitudes of displacement vector between point pairs. Scenarios (a) and (b) of Figure 6 show the effectiveness of Euclidean Relation Clues and more discussion can be viewed in § 4.4.
@@ 3.4 Euclidean Loss and Objectives
We transform both the ground truth data and the model’s predicted polylines from their original format into G-Representation. This conversion allows the model to independently optimize the angles between displacement vectors and the magnitudes of displacement vectors to more accurately align with the ground truth. The proposed Euclidean Loss is composed of two parts that measure how accurately the model predicts the shape of individual map instances and the inter-instance relations, respectively, which are denoted by $\mathcal{L}_{\rm shp}$ and $\mathcal{L}_{\rm rel}$. The Euclidean Loss can be computed as the following:
$\mathcal{L}_{\rm Euc}=\lambda_{1}\cdot\mathcal{L}_{\rm shp}+\lambda_{2}\cdot\mathcal{L}_{\rm rel},$ (4)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters, whose effects on the model performances are evaluated in Figure LABEL:fig:shape_rel. Specifically,
[…]
$\mathcal{L}_{\rm rel}$ $=\sum_{i=1}^{N}\sum_{j>i}^{N}w_{i,j}\sum_{u=1}^{N_{v}}\sum_{v=1}^{N_{v}}\left(|\hat{d}^{i,j}_{u,v}-d^{i,j}_{u,v}|+\ell(\hat{\alpha}^{i,j}_{u,v},\alpha^{i,j}_{u,v})\right)\,,$
where $\ell(\cdot)$ is a function based on $L^{1}$ loss. We avoid using the inverse trigonometric function to directly compute the angles, but use sine and cosine values instead:
@@
Moreover, the distance might influence the relation strength of instance pairs. For example, if two instances are far from each other, the relation between them might be weak. For this reason, we further adopt $w_{i,j}$ to punish weakly related instance pairs and discuss it in § 4.5. According to the best experimental results, we treat all instance pairs equally in other experiments.
Following the common practice [20, 46], we also use focal loss [22] $\mathcal{L}_{\rm cls}$ for classification. For polyline regression, we adopt point-to-point loss and edge direction loss, denoted by $\mathcal{L}_{\rm pts}$ and $\mathcal{L}_{\rm dir}$ respectively. Following [34, 21], we further adopt segmentation loss $\mathcal{L}_{\rm seg}$ and depth estimation loss $\mathcal{L}_{\rm dep}$. These losses are detailed in appendix § 0.A.1. The overall loss function can be written as:
$\mathcal{L}$ $=\lambda\cdot\mathcal{L}_{\rm Euc}+\beta_{1}\cdot\mathcal{L}_{\rm cls}+\beta_{2}\cdot\mathcal{L}_{\rm pts}+\beta_{3}\cdot\mathcal{L}_{\rm dir}+\beta_{4}\cdot\mathcal{L}_{\rm seg}+\beta_{5}\cdot\mathcal{L}_{\rm dep}.$ (7)
where hyperparameters are discussed in Figure LABEL:fig:lossw and appendix Table A1.
3.5 Geometry-Decoupled Decoder
In the Transformer decoder, we obtain the $N\times N_{v}$ fused queries by adding $N$ instance queries to $N_{v}$ point queries, which is suggested in [20]. We denote input tokens by ${\bm{E}}\in\mathbb{R}^{N_{A}\times D}$, where $D$ is the feature dimension and $N_{A}=N\cdot N_{v}$ is the number of fused queries. Intuitively, each fused query corresponds to a point on a predicted map instance. Self-Attention (SA) is formulated by:
${\rm SA}({\bm{E}},{\bm{M}})={\rm Softmax}\left(\frac{({\bm{E}}{\bm{W}}^{q})({\bm{E}}{\bm{W}}^{k})^{\top}}{\sqrt{D_{k}}}\odot{\bm{M}}\right){\bm{E}}{\bm{W}}^{v}\,,$ (8)
Therefore, we present Geometry-Decoupled Attention (GDA). First, we multiply a binary mask ${\bm{M}}$ to the computed attention map so that the tokens of a map instance are aggregated according to the tokens within the same map instances only, which allows the model to adjust points’ positions according to learned shape geometry. Denote the index of map instance that the $i$-th token belongs to as $\mathcal{I}_{i}$. For example, tokens 1,2$\dots$,$N_{v}$ belong to the first instance so that $\mathcal{I}_{1},\mathcal{I}_{2},\dots,\mathcal{I}_{N_{v}}=1$. Tokens $N_{v}$+1, $N_{v}$+2, $\dots$, $2\cdot N_{v}$ belong to the second instance so that $\mathcal{I}_{N_{v}+1},\mathcal{I}_{N_{v}+2},\dots,\mathcal{I}_{2N_{v}}=2$. With this notation, the binary mask ${\bm{M}}$ can be simply constructed by:
The second attention is expected to model relations between tokens of different map instances. The attention mask is given by:
4 Experiments
4.1 Experimental Setups
Table 1: Comparison on the nuScenes dataset, GeMap reaches a new state-of-the-art performance. “EB0”, “R50”, “PP”, “Sec”, “Swin-T”, and “V2-99” denote EfficientNet-B0 [38], ResNet50 [10], PointPillars [13], SECOND [44], Swin Transformer Tiny [25], and VoVNetV2-99 [14] respectively. Methods with two backbones utilize both camera and LiDAR inputs. “Dns. Loss” denotes whether any dense prediction (e.g. semantic segmentation) loss is adopted. The best result is highlighted in bold. We reproduce all methods on a single RTX3090 GPU to test FPS for fair comparison.
Methods Backbone Dns. Loss APdiv()($\uparrow$)( ↑ ) APped()($\uparrow$)( ↑ ) APbnd()($\uparrow$)( ↑ ) mAP()($\uparrow$)( ↑ ) FPS()($\uparrow$)( ↑ )